OpenStudy (anonymous):

Find the relation between p,q and r if the roots of the cubic equation x^3 -px^2+qx -r =0 are such that they are in AP

ganeshie8 (ganeshie8):

say the roots are \(\alpha, ~\alpha + d,~ \alpha + 2d\)

ganeshie8 (ganeshie8):

sum of roots = p sum of product of pairs of roots = q product of roots = r

OpenStudy (anonymous):

ok....

ganeshie8 (ganeshie8):

we get 3 equations from above^

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

we need to eliminate \(\alpha, ~ d\) from them and get a relation in terms of p, q, r i guess..

OpenStudy (anonymous):

wait let me type the equations

ganeshie8 (ganeshie8):

k

OpenStudy (dan815):

the roots of a cubic equation vary with constant difference?

OpenStudy (dan815):

a,a+d,a+2d? is this always true

OpenStudy (dan815):

if 3 roots exist

OpenStudy (anonymous):